OpenStudy (anonymous):

f(x)= 7x/(x^2-9). use interval notation to indicate where f(x) is decreasing. what is the local min and max and inflection points

OpenStudy (anonymous):

First find the derivative \[ f'(x)=\frac {7(x^2-9) - 2x(7x) } {(x^2 -9)^2}= \frac {-7 x^2-63 } {(x^2 -9)^2}=-\frac {7 (x^2+9) } {(x^2 -9)^2} \]

OpenStudy (anonymous):

What can you say about the sign of the derivative?

OpenStudy (anonymous):

the negative sign? I'm not sure what do you mean

OpenStudy (anonymous):

Is \[x^2 + 9 \] always >0?

OpenStudy (anonymous):

no... you can't take the root of a negative number!

OpenStudy (anonymous):

Is \[ (x^2 -9)^2 \ge 0 \]

OpenStudy (anonymous):

Give me a real number x such that \[ x^2 + 9 \le 0 \] There is not such a number. \[ x^2 \ge 0\\ 9 > 0\\ x^2 + 9 > 0 \]

OpenStudy (anonymous):

so its none? or does not exist?

OpenStudy (anonymous):

I mean \[ x^2 + 9\] is always strictly bigger than zero.

OpenStudy (anonymous):

how do i find the interval that it is decreasing?

OpenStudy (anonymous):

The derivative \[ -\frac {7 (x^2+9) } {(x^2 -9)^2} \] is always < 0 when it is defined. So f is decreasing everywhere in its domain. The domain is every x except x=-3 and x =3.

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

ya thanks!

OpenStudy (anonymous):

so to find out if it is concave up or down i take the derivative again right?

OpenStudy (anonymous):

Can this function have local max or local min?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

no it can not

OpenStudy (anonymous):

great.

OpenStudy (anonymous):

I'm supposed to find interval notation of when it is concave up and when it is concave down. so the second derivative numerator is 14x(x^2+27)... so x^2+27>0? i don't get what i do after to find the intervals

OpenStudy (anonymous):

Now try to find the second derivative step by step. To check your answer, \[ f''(x)=\frac{14 x \left(x^2+27 \right)}{\left(x^2-9\right )^3} \] You can argue that the sign of f''(x) is the sign of 14 x. Why?

OpenStudy (anonymous):

because then you distribute it? do you mean why it is positive?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that x is greater than zero since our f(x) is positive?

OpenStudy (anonymous):

No, that means that x=0 is an inflection point.

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

There are two other points x=-3 and x=3 where f has vertical asymptotes where the concavity changes.

OpenStudy (anonymous):

The three points that you should worry about are x -Infinity -3 0 3 +Infinity f'(x) - + 0 - + f(x) conc dn con up conc dn conc up

OpenStudy (anonymous):

f''(x) in the second rwo

OpenStudy (anonymous):

ohhh i get it! perfect

OpenStudy (anonymous):

f is concave up on (-3,0) and (3, Infinity) Where is concave down?

OpenStudy (anonymous):

(-inf, -3) and (0,3)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you!!!!!!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

:)